# Parameter estimation of microlensed gravitational waves with Conditional Variational Autoencoders

### v1

@@ Abstract
Gravitational lensing of gravitational waves (GWs) provides a unique opportunity to study cosmology and astrophysics at multiple scales. Detecting microlensing signatures, in particular, requires efficient parameter estimation methods due to the high computational cost of traditional Bayesian inference. In this paper we explore the use of deep learning, namely Conditional Variational Autoencoders (CVAE), to estimate parameters of microlensed binary black hole (simulated) waveforms. We find that our CVAE model yields accurate parameter estimation and significant computational savings compared to Bayesian methods such as bilby (up to five orders of magnitude faster inferences). Moreover, the incorporation of CVAE-generated priors in bilby reduces the average runtime of the latter in about 48% with no penalty on its accuracy. Our results suggest that a CVAE model is a promising tool for future low-latency searches of lensed signals. Further applications to actual signals and integration with advanced pipelines could help extend the capabilities of GW observatories in detecting microlensing events.
@@ I Introduction
The first three observing runs of the network of gravitational-wave (GW) detectors Advanced LIGO [1], Advanced Virgo [2] and KAGRA [3] have led to the detection of 90 compact binary coalescence (CBC) events [4, 5, 6, 7]. Their analysis is impacting different fields of research, e.g. astronomy, astrophysics, cosmology, and fundamental physics, in a significant way. In particular the wealth of new data has provided updated information on the CBC merger rate and has allowed for new studies of populations of compact objects (specifically regarding the distribution of black hole masses) and for new tests of general relativity [8, 9]. One distinct aspect being scrutinized is the possibility of finding evidence of gravitational lensing on the observed GW signals [10, 11]. If it took place, lensing by massive objects along the line of sight between the detector and the source might significantly distort the signal leading to e.g. beating patterns, signal amplitude changes, signal repetition, or a combination of these phenomena.
@@
The LIGO-Virgo-KAGRA (LVK) Collaboration has conducted several searches for GW lensing, including both strong lensing and microlensing signatures. The most recent searches have accounted for all signals from the whole sample of binary black hole (BBH) mergers reported in the third observing run [10, 11]. None of the searches conducted in those studies has yielded conclusive evidence for GW lensing [12]. Other groups, however, have put forward different proposals suggesting that the high masses observed in some BBH systems could be explained by the magnification effect of gravitational lensing [13, 14, 15] (see also related discussions in [10, 16]).
@@
Despite lensing of gravitational waves has not yet been established, it is firmly regarded as a new instrument for GW astrophysics with a remarkable potential for new discoveries, as highlighted in studies related to the LISA mission [17]. Certain gravitational lenses might be composed of objects that neither emit light nor neutrinos, making them undetectable through conventional means. Observing GW lensing could hence provide crucial insights into populations of elusive or poorly understood objects, such as intermediate-mass black holes [18, 19, 20], dark matter haloes [21, 22, 23], cosmic strings [24, 25, 26], or even potentially reveal entirely new classes of objects. The first detection of gravitational lensing by such phenomena would provide direct evidence of compact dark matter objects, offering a groundbreaking advance in astrophysics [27, 28, 21, 29, 30, 22, 31, 32, 33, 34, 35, 36].
Just as it happens for electromagnetic signals, the effects and impact of lensing on a GW can be remarkably diverse. On the one hand, lensing amplification can cause the source’s distance and mass to be underestimated, leading to incorrect estimation of parameters. If a GW is lensed, it might appear closer and more massive than it truly is, which makes the identification of the effect crucial to avoid such errors. On the other hand, strong lensing, caused by galaxies, would produce repeated GW events separated by minutes to months. This phenomenon occurs when large masses significantly bend spacetime, creating multiple images of the source [37, 38, 39]. In contrast, microlensing, caused by stars or compact objects, results in much shorter time delays, with duplicate events separated by milliseconds to seconds. Unlike strong lensing, microlensing can magnify signals or create a “beating” pattern in the waveform, where the time delay between paths is comparable to the wave’s period.
The investigation of the lensing phenomenon has received significant attention in GW astronomy, with an increasing body of work that offers unique insights into small-scale structures in the universe [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. In this paper we report on our contribution to these efforts by discussing the results from a study on the microlensing case, particularly in the point-mass lens model. This is the simplest and most illustrative model for a microlensing object, where the lensing effects depend only on the redshifted lens mass $M_{Lz}$ and on the parameter $y$, which defines the position of the source relative to the line of sight.
@@
The straightforward approach to detecting microlensing is to do Bayesian parameter estimation of individual events with lensing added to the waveform model. Currently, the pipeline in charge of performing these searches is Gravelamps [59]. The biggest limitation of a Bayesian approach is computational efficiency, as it can be extremeley time consuming. Recent studies [75, 76] have demonstrated that deep-learning techniques, particularly Conditional Variational Autoencoders (CVAE) [77], offer a powerful way to perform fast parameter estimation in (unlensed) GW data. Indeed, CVAEs have shown impressive potential in capturing complex features of GW signals, making them a valuable tool in contexts where speed is essential (e.g. low-latency searches). Other deep-learning approaches have been used for the identification of microlensed [78] and strong-lensed [79, 80] signals, as well as on parameter estimation and detection of strong lensing in the electromagnetic case [81, 82]. Despite these advances, there are currently no studies that apply deep-learning techniques specifically to the case of parameter estimation of microlensed gravitational waves. The aim of the present investigation is to bridge that gap, leveraging the unique capabilities of CVAEs to address the complex problem of parameter estimation in microlensing scenarios.
@@ II Microlensing
Any astrophysical object with sufficient mass can act as a lens for a passing gravitational wave. Lensing effects occur when the source, the lens and the observer are all aligned within the Einstein angle $\theta_{\rm E}=R_{\rm E}/d_{L}$, i.e., the lens is located near the line of sight. The Einstein radius $R_{\rm E}$ is commonly expressed as:
@@
where $M_{L}$ is the mass of the lens, $d_{LS}$ is the angular diameter distance between the source and the lens, and $d_{L}$ and $d_{S}$ are the angular diameter distances to the lens and source at redshifts $z_{L}$ and $z_{S}$, respectively. $R_{\rm E}$ represents the characteristic length scale on the lens plane, is proportional to $\sqrt{M_{L}}$ and is typically much smaller than the cosmological distances $d_{LS}$, $d_{L}$, and $d_{S}$. This enables the lens mass to be projected onto a lens plane. In the thin-lens approximation, gravitational waves propagate freely outside the lens, interacting only with a two-dimensional gravitational potential at the lens plane, where the lensing effect is ultimately captured in the transmission factor $F$ [84].
@@
where $\nu\equiv ft_{M}$ is the frequency of the gravitational wave $f$ scaled by the characteristic time $t_{M}$, $\Gamma(z)$ is the Gamma function, and ${}_{1}F_{1}(a,b,z)$ is the confluent hypergeometric function [^1] In our code we take the complex conjugate by replacing $\mathrm{i}$ with $-\mathrm{i}$ to align with the Fourier transform convention used in the Python libraries of the LVK Collaboration [88].. With the chosen normalization $\xi_{0}=R_{\rm E}$, the time $t_{M}$, as expressed by Eq. (5), is proportional to the mass of the lens:
@@
We are interested in the behavior of the transmission factor within the frequency range detectable by the LVK network, which spans approximately from 30 Hz to 1 kHz [89, 90, 91, 92, 93, 94]. For a given choice of parameters $M_{Lz}$ and $y$, the lensing effect on GWs in the LVK range can fall into the amplification region, the oscillating region, or an intermediate region, depending on the lens mass [61]. For lower masses (e.g. $M_{L}=30M_{\odot}$), the primary effect is amplification. As the mass increases, oscillatory effects with a beating pattern emerge in the waveform [65]. At high frequencies, calculating the transmission factor (8) in the oscillating region becomes numerically expensive. Fortunately, in this limit, the dominant contribution arises from two well-defined images of the source, corresponding to two stationary points of the time delay function. The transmission factor in this geometrical optics (GO) limit can then be written as a sum over these stationary points [84, 41, 42],
@@ III Conditional Variational autoencoder
where $p(h|\Lambda)$ represents the likelihood and $p(\Lambda)$ the prior of the parameters. The common way of approaching this problem with Bayesian inference methods involves several explicit likelihood evaluations, which results in large computational times [95]. In this context, likelihood-free, machine learning methods have been proposed as a way of speeding up the process [75, 76]. A kind of models that have proved specially useful for the estimation of posterior probability distributions are CVAEs. Autoencoders [96] are a type of neural network used to learn efficient, compressed representations of data in a space – the so-called latent space – of lower dimension than the input. An autoencoder consists of two main parts:
@@
In this work, we construct a CVAE following the lines drawn in [75], specializing it for the estimation of the two parameters determining the lensing effect of a PML, $\Lambda_{L}=\{t_{M},y\}$. This CVAE consists on a pair encoder-decoder, $r_{\theta_{1}}$ and $r_{\theta_{2}}$, together with a recognition encoder, $q_{\phi}(z|\Lambda,h)$, where $\theta_{1}$, $\theta_{2}$ and $\phi$ represent the set of trainable parameters of each network and $\Lambda=\{d_{L},m_{1},m_{2},t_{M},y\}$ are the parameters fed to the recognition encoder. The small changes done to the network with respect to the one in [75] are detailed in Appendix A, as well as the details concerning the training process. A scheme of the network used for the training phase can be seen in Fig. 1. Correspondingly, a scheme of the test phase is shown in Fig. 2. Notice that the part of conditioning information mentioned in the last paragraph, is played in this model by the GW strain in the time domain, $h(t)$.
@@ IV Data
Figure 3: Example of a whitened, lensed BBH waveform from the datasets. The waveform is shown in red and the noise in blue. In this case, lensing is strong enough to generate two images.
In this study we rely on utilities from the bilby software package [83] in order to generate both lensed and unlensed BBH merger waveforms. The process of generation is as follows: first, we generate noise data for the Hanford (H1) and Livingston (L1) LIGO detectors at a sampling rate of 2048 Hz, following the methodology described in [100]. Specifically, we use Dataset 3, which consists of purely Gaussian noise, sampled according to power spectral densities calculated from real data from the O3a observation run. Then, we sample the physical parameters of a given waveform from the default bilby BBH prior values save for the component masses, which are uniformly sampled in the $[35,80]\leavevmode\nobreak\ M_{\odot}$ interval and forced to satisfy $m_{1}>m_{2}$. We additionally sample the lensing parameters $y\in\left[0.01,3.0\right]$ and $t_{M}\in[0.001,0.2]$ seconds, as well as a target signal-to-noise ratio (SNR) in the interval $\textrm{SNR}\in[8,40]$. The priors for the rest of parameters are the default priors used by bilby. The physical parameters are then passed to the waveform generator, which uses the IMRPhenomXPHM approximant [101]. This waveform generator computes the transmision factor following the approach explained in Section II. Lensing is applied to the resulting frequency-domain waveform according to
@@
The sizes of the training, validation and testing datasets are, respectively, $10^{6}$, $2\times 10^{4}$ and $250$. These datasets do not intersect to ensure that we can monitorize the generalization ability of the model. An example of a whitened, lensed waveform from one of the datasets is shown in Fig. 3.
@@ V Results
We train our CVAE using the architecture detailed in Appendix A, achieving a smooth evolution of the loss function, as shown in Fig. 4 (specific definitions are provided in the appendix). During the initial 100 epochs, while the annealing process is active, the loss shows greater fluctuations. After this period, when annealing ends and the learning rate is set to $10^{-5}$, the loss evolution becomes noticeably smoother. The loss converges to approximately -2.5, with no signs of overfitting, as both training and validation losses remain close throughout.
@@
At the end of training, we save the model from the epoch with the best performance, specifically the epoch with the lowest validation loss, and use it for testing. Using the trained CVAE, we generate 8,000 samples for each of the 250 test waveforms, produced following the guidelines defined in Section IV. These samples define the posterior distributions for each parameter, $y$ and $t_{M}$. A probability-probability (p-p) plot comparing the distribution of parameters of the test set with the distribution of the generated posterior samples is shown in Figure 5. A p-p plot compares the cumulative distribution of predicted probabilities to the cumulative distribution of observed data. In this plot, the $x$-axis represents the quantiles of the cumulative probability predicted by the model (e.g. the quantile range 0% to 100%), whereas the $y$-axis represents the observed frequency of true values falling within these predicted quantile ranges. When the plot lines are close to the diagonal line, it indicates that the model’s predicted probability intervals match the observed data well. In other words, for an $X\%$ quantile on the $x$-axis, approximately $X\%$ of the true values fall within that quantile range in the observed data. A p-p plot close to the diagonal line suggests that the model is well-calibrated: the predicted uncertainty aligns well with the actual uncertainty in the data. As Fig. 5 shows, our trained CVAE yields satisfactory results for both model parameters.
@@
The scatter plots in Figure 7 show the test set waveforms, labeled by their values of $y$ and $t_{M}$. The color code in the dots indicates whether the waveforms are correctly (blue) or incorrectly (red) processed by the CVAE. Notably, the CVAE successfully estimates $y$ and $t_{M}$ even when $t_{M}$ is close to 0 (i.e. when the lensing effect is nearly absent), demonstrating its ability to handle unlensed waveforms. Additionally, the model accurately detects lensing for values of $y>1$, where the effect is weaker. These cases are challenging to identify, and previous work, such as [78], did not even account for waveforms with those $y$ values.
@@
Finally, we analyze the efficiency of the CVAE to perform the inference of the lens parameters, comparing our timing estimates to those of bilby. In order to do so we perform parameter estimation on 17 different BBH waveforms and measure the inference (wall-clock) time [^2] The computations were carried out at the ICCUB Nyx’s cluster, equipped with dual AMD EPYC 7763 processors, each with 64 cores.. The results are presented in Table 1. The values reported in this table are obtained from those of Table 3 in Appendix A which shows all timing estimates for the entire set of waveforms. We observe that the CVAE is significantly more efficient than Bilby, achieving speeds about 4 orders of magnitude faster (occasionally even 5 orders of magnitude; see Table 3). To accelerate Bilby’s convergence, we use the true values as priors for all parameters except $y$ and $t_{M}$. For $y$ and $t_{M}$, we set uniform priors ranging from 0 to 3 and from 0 to 0.2 seconds, respectively. When the CVAE is used to provide priors for Bilby, the priors for these two lensing parameters are uniform distributions defined by the 95% confidence intervals of the CVAE’s posterior estimates, i.e. $\mu\pm 2\sigma$, where $\mu$ and $\sigma$ represent the posterior mean and standard deviation respectivelt. In this setup, the average runtime is reduced by 47.9% compared to the previous configuration.
@@
In addtion to significantly speed up the inference, the use of the CVAE, either combining it with bilby or otherwise, does not degrade the accuracy of the parameter estimation obtained by bilby when used in standalone mode. A proof of this assertion is shown in Figure 8 which depicts a corner plot of the posterior distributions of the two lens parameters, for a particular waveform of our sample, obtained with both bilby and the CVAE in standalone mode, plus the corresponding estimation made with the two approaches combined.
Figure 8: Example of parameter estimation with bilby in standalone mode (blue), the CVAE (red) and a combination of both (green) for the same waveform. Notice that the blue and green posteriors practically overlap, showing that providing priors with CVAE does not degrade the accuracy of the estimations made by bilby.

### v2

@@ Abstract
Gravitational lensing of gravitational waves (GWs) provides a unique opportunity to study cosmology and astrophysics at multiple scales. Detecting microlensing signatures, in particular, requires efficient parameter estimation methods due to the high computational cost of traditional Bayesian inference. In this paper we explore the use of deep learning, namely Conditional Variational Autoencoders (CVAE), to estimate parameters of microlensed binary black hole (simulated) waveforms. We find that our CVAE model yields accurate parameter estimation and significant computational savings compared to Bayesian methods such as bilby (up to five orders of magnitude faster inferences). Moreover, the incorporation of CVAE-generated priors into bilby, based on the 95% confidence intervals of the CVAE posterior for the lensing parameters, reduces Bilby’s average runtime by around 48% without any penalty on accuracy. Our results suggest that a CVAE model is a promising tool for future low-latency searches of lensed signals. Further applications to actual signals and integration with advanced pipelines could help extend the capabilities of GW observatories in detecting microlensing events.
@@ I Introduction
The first three observing runs of the network of GW detectors Advanced LIGO [1], Advanced Virgo [2] and KAGRA [3] have led to the detection of 90 compact binary coalescence (CBC) events [4, 5, 6, 7]. Their analysis is impacting different fields of research, e.g. astronomy, astrophysics, cosmology, and fundamental physics, in a significant way. In particular the wealth of new data has provided updated information on the CBC merger rate and has allowed for new studies of populations of compact objects (specifically regarding the distribution of black hole masses) and for new tests of general relativity [8, 9]. One distinct aspect being scrutinized is the possibility of finding evidence of gravitational lensing on the observed GW signals [10, 11]. If it took place, lensing by massive objects along the line of sight between the detector and the source might significantly distort the signal leading to e.g. beating patterns, signal amplitude changes, signal repetition, or a combination of these phenomena.
@@
The LIGO-Virgo-KAGRA (LVK) Collaboration has conducted several searches for GW lensing, including both strong lensing and microlensing signatures. The most recent searches have accounted for all signals from the whole sample of binary black hole (BBH) mergers reported in the third observing run [10, 11]. None of the searches conducted in those studies has yielded conclusive evidence for GW lensing [11, 12]. Other groups, however, have put forward different proposals suggesting that the high masses observed in some BBH systems could be explained by the magnification effect of gravitational lensing [13, 14, 15] (see also related discussions in [10, 16]).
@@
Although GW lensing has yet to be conclusively observed, it is widely recognized as a promising tool for GW astrophysics, offering significant potential for groundbreaking discoveries. This potential has been underscored in studies involving both third-generation terrestrial GW observatories [17, 18] and the space-based LISA mission [19]. Certain gravitational lenses might be composed of objects that neither emit light nor neutrinos, making them undetectable through conventional means. Observing GW lensing could hence provide crucial insights into populations of elusive or poorly understood objects, such as intermediate-mass black holes [20, 21, 22], dark matter haloes [23, 24, 25], cosmic strings [26, 27, 28], or even potentially reveal entirely new classes of objects. The first detection of gravitational lensing by such phenomena would provide direct evidence of compact dark matter objects, offering a groundbreaking advance in astrophysics [29, 30, 23, 31, 32, 24, 33, 34, 35, 36, 37, 38].
Lensing effects on GWs, much like their electromagnetic counterparts, can manifest in a variety of regimes, each with distinct characteristics and implications. In the strong lensing regime, massive structures such as galaxies curve spacetime enough to create multiple images of the source. These repeated GW events are typically separated by time delays ranging from minutes to months [39, 40, 41]. In contrast, the microlensing regime arises from smaller-scale objects like stars or stellar-mass compact objects, where the time delays are much shorter, typically spanning milliseconds to seconds [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. This regime can amplify signals or imprint a distinctive ’beating’ pattern in the waveform, caused by time delays that are comparable to the wave’s period111 Some authors further distinguish between microlensed GW signals, occurring in the wave optics regime, and millilensed GW signals, which fall within the geometric optics regime and are associated with comparatively higher lens masses [68, 69]. Given that our research covers a broad range of masses encompassing both the wave-optics and geometric-optics limits (utilizing the hybrid transmission factor as detailed later), we will collectively refer to both regimes as ’microlensing’ without distinction, consistent with the classification adopted in Refs. [10, 11].. Microlensing has profound astrophysical implications and has garnered considerable attention in GW astronomy, as it provides unique insights into small-scale structures in the universe. For example, lensing amplification can bias the inferred parameters of a source, such as its distance and mass, if left unaccounted for. A lensed GW may appear closer and more massive than it truly is, underscoring the importance of identifying lensing effects to avoid systematic errors.
In this paper, we contribute to these efforts by focusing on the microlensing regime and presenting results from a study based on the point-mass lens model. This model, the simplest and most illustrative for microlensing, captures lensing effects as a function of the redshifted lens mass $M_{Lz}$ and the source position parameter $y$, defined relative to the line of sight.
@@
The straightforward approach to detecting microlensing is to do Bayesian parameter estimation of individual events with lensing added to the waveform model. Currently, the pipeline in charge of performing these searches is Gravelamps [61]. The biggest limitation of a Bayesian approach is computational efficiency, as it can be extremeley time consuming. Recent studies have demonstrated that deep-learning techniques offer a powerful way to perform fast parameter estimation in (unlensed) GW data (see the comprehensive review [80] and refrences therein). In particular, Conditional Variational Autoencoders (CVAE) [81, 82], normalizing flows [83, 84], together with its extensions Nessai [85, 86], Dingo [87, 88] and flow matching [89], have shown impressive potential in capturing complex features of GW signals, making them a valuable tool in contexts where speed is essential (e.g. low-latency searches). Notably, CVAEs have already demonstrated their versatility in addressing a variety of GW analysis challenges, including continuous waves [90], ringdown signals [91], and postmerger signals from binary neutron stars [92], among others. Despite these advances, neither CVAEs nor normalizing flows have yet been applied to the parameter estimation of lensed GWs. Other deep-learning approaches have been used for the identification of microlensed [93] and strong-lensed [94, 95] signals, as well as on parameter estimation and detection of strong lensing in the electromagnetic case [96, 97]. The aim of the present investigation is to bridge that gap, leveraging the unique capabilities of CVAEs to address the complex problem of parameter estimation in microlensing scenarios.
@@ II Microlensing
Any astrophysical object with sufficient mass can act as a lens for a passing GW. Lensing effects occur when the source, the lens and the observer are all aligned within the Einstein angle $\theta_{\rm E}=R_{\rm E}/d_{L}$, i.e., the lens is located near the line of sight. The Einstein radius $R_{\rm E}$ is commonly expressed as:
@@
where $M_{L}$ is the mass of the lens, $d_{LS}$ is the angular diameter distance between the source and the lens, and $d_{L}$ and $d_{S}$ are the angular diameter distances to the lens and source at redshifts $z_{L}$ and $z_{S}$, respectively. $R_{\rm E}$ represents the characteristic length scale on the lens plane, is proportional to $\sqrt{M_{L}}$ and is typically much smaller than the cosmological distances $d_{LS}$, $d_{L}$, and $d_{S}$. This enables the lens mass to be projected onto a lens plane. In the thin-lens approximation, GWs propagate freely outside the lens, interacting only with a two-dimensional gravitational potential at the lens plane, where the lensing effect is ultimately captured in the transmission factor $F$ [99].
@@
where $\nu\equiv ft_{M}$ is the frequency of the GW $f$ scaled by the characteristic time $t_{M}$, $\Gamma(z)$ is the Gamma function, and ${}_{1}F_{1}(a,b,z)$ is the confluent hypergeometric function [^2] In our code we take the complex conjugate by replacing $\mathrm{i}$ with $-\mathrm{i}$ to align with the Fourier transform convention used in the Python libraries of the LVK Collaboration [103].. With the chosen normalization $\xi_{0}=R_{\rm E}$, the time $t_{M}$, as expressed by Eq. (5), is proportional to the mass of the lens:
@@
We are interested in the behavior of the transmission factor within the frequency range detectable by the LVK network, which spans approximately from 30 Hz to 1 kHz [104, 105, 106, 107, 108, 109]. For a given choice of parameters $M_{Lz}$ and $y$, the lensing effect on GWs in the LVK range can fall into the amplification region, the oscillating region, or an intermediate region, depending on the lens mass [63]. For lower masses (e.g. $M_{L}=30M_{\odot}$), the primary effect is amplification. As the mass increases, oscillatory effects with a beating pattern emerge in the waveform [67]. At high frequencies, computing the transmission factor (8) in the oscillating region becomes numerically expensive. Fortunately, in this limit, the dominant contribution arises from two well-defined images of the source, corresponding to two stationary points of the time delay function. In general, the transmission factor in this geometrical optics limit can be written as a sum over all stationary points [99, 43, 44],
@@ III Conditional Variational autoencoder
where $p(h|\Lambda)$ represents the likelihood and $p(\Lambda)$ the prior of the parameters. The common way of approaching this problem with Bayesian inference methods involves several explicit likelihood evaluations, which results in large computational times [110]. In this context, likelihood-free [111] machine learning methods have been proposed as a way of speeding up the process [81, 82, 83, 84, 85, 86, 87, 88, 89].
A kind of models that have proved specially useful for the estimation of posterior probability distributions are CVAEs. Autoencoders [112] are a type of neural network used to learn efficient, compressed representations of data in a space – the so-called latent space – of lower dimension than the input. An autoencoder consists of two main parts:
@@
In this work, we construct a CVAE following the lines drawn in [82], specializing it for the estimation of the two parameters determining the lensing effect of a PML, $\Lambda_{L}=\{t_{M},y\}$. This CVAE consists on a pair encoder-decoder, $r_{\theta_{1}}$ and $r_{\theta_{2}}$, together with a recognition encoder, $q_{\phi}(z|\Lambda,h)$, where $\theta_{1}$, $\theta_{2}$ and $\phi$ represent the set of trainable parameters of each network and $\Lambda=\{d_{L},m_{1},m_{2},t_{M},y\}$ are the parameters fed to the recognition encoder. The other GW parameters were not fed to the model in an effort to reduce the number of trainable parameters of the CVAE. The small changes done to the network with respect to the one in [82] are detailed in Appendix A, as well as the details concerning the training process. A scheme of the network used for the training phase can be seen in Fig. 1. Correspondingly, a scheme of the test phase is shown in Fig. 2. Notice that the conditioning information given to the CVAE is precisely the time series of the GW, $h(t)$.
@@ IV Data
In this study we rely on utilities from the Bilby software package [98] in order to generate both lensed and unlensed BBH merger waveforms. The process of generation is as follows: first, we generate noise data for the Hanford (H1) and Livingston (L1) LIGO detectors at a sampling rate of 2048 Hz, following the methodology described in [116]. Specifically, we use Dataset 3, which consists of purely Gaussian noise, sampled according to power spectral densities calculated from real data from the O3a observation run. Then, we sample the physical parameters of a given waveform from the default Bilby BBH prior values save for the component masses, which are uniformly sampled in the $[35,80]\leavevmode\nobreak\ M_{\odot}$ interval and forced to satisfy $m_{1}>m_{2}$, following [82]. We additionally sample the lensing parameters $y\in\left[0.01,3.0\right]$ and $t_{M}\in[0.001,0.2]$ seconds, as well as a target signal-to-noise ratio (SNR) in the interval $\textrm{SNR}\in[8,40]$. The specified range of $t_{M}$ corresponds to lens masses within $50<M_{Lz}/M_{\odot}<10^{4}$. The priors for the rest of parameters are the default priors used by Bilby. The physical parameters are then passed to the waveform generator, which uses the IMRPhenomXPHM approximant [117]. This waveform generator computes the transmision factor following the approach explained in Section II. Lensing is applied to the resulting frequency-domain waveform according to
@@
The sizes of the training, validation and testing datasets are, respectively, $10^{6}$, $2\times 10^{4}$ and 1000. These datasets do not intersect to ensure that we can monitorize the generalization ability of the model. An example of a whitened, lensed waveform from one of the datasets is shown in Fig. 3.
Figure 3: Example of a whitened, lensed BBH waveform from the datasets. The signal is depicted in red, while its noise-masked version appears in blue. For this signal, with parameters: $t_{M}=0.121\,$s, $y=0.842$, the lensing effect is strong enough to produce two distinct images with an SNR of $25.3$.
@@ V Results
We train our CVAE using the architecture detailed in Appendix A, achieving a smooth evolution of the loss function, as shown in Fig. 4 (specific definitions are provided in the appendix). During the initial 100 epochs, while the annealing process is active, the loss shows greater fluctuations. After this period, when annealing ends and the learning rate is set to $10^{-5}$, the loss evolution becomes noticeably smoother. The loss converges to approximately -2.5, with no signs of overfitting, as both training and validation losses remain close throughout. The training process took between 7 and 8 hours.
@@
At the end of training, we save the model from the epoch with the best performance—specifically, the epoch with the lowest validation loss, which occurred at epoch 1078—and use it for testing. Using the trained CVAE, we generate 8000 samples for each of the 1000 test waveforms, produced following the guidelines defined in Section IV. These samples define the posterior distributions for the two parameters estimated by the CVAE, $y$ and $t_{M}$. A probability-probability (p-p) plot comparing the distribution of parameters of the test set with the distribution of the generated posterior samples is shown in Figure 5. A p-p plot compares the cumulative distribution of predicted probabilities to the cumulative distribution of observed data. In this plot, the $x$-axis represents the quantiles of the cumulative probability predicted by the model (e.g. the quantile range 0% to 100%), whereas the $y$-axis represents the observed frequency of true values falling within these predicted quantile ranges. When the plot lines are close to the diagonal line, it indicates that the model’s predicted probability intervals match the observed data well. In other words, for an $X\%$ quantile on the $x$-axis, approximately $X\%$ of the true values fall within that quantile range in the observed data. A p-p plot close to the diagonal line suggests that the model is well-calibrated: the predicted uncertainty aligns well with the actual uncertainty in the data. As Fig. 5 shows, our trained CVAE yields satisfactory results for both model parameters.
@@
The scatter plots in Figure 7 show the test set waveforms, labeled by their values of $y$ and $t_{M}$. The color code in the dots indicates whether the waveforms are correctly (blue) or incorrectly (red) processed by the CVAE, i.e., if the true values lie within the 95% credible intervals defined by the posterior generated by the CVAE. Notably, the CVAE successfully estimates $y$ and $t_{M}$ even when $t_{M}$ is close to 0 (i.e. when the lensing effect is nearly absent), demonstrating its ability to handle unlensed waveforms. Additionally, the model accurately estimates the two lensing parameters for values of $y>1$, where the effect is weaker. These cases are challenging to identify, and previous work, such as [93], did not even account for waveforms with those $y$ values.
@@
Finally, we analyze the efficiency of the CVAE to perform the inference of the lens parameters, comparing our timing estimates to those of Bilby. In order to do so we perform parameter estimation on 17 different BBH waveforms and measure the inference (wall-clock) time [^3] The computations were carried out at the ICCUB Nyx’s cluster, equipped with dual AMD EPYC 7763 processors, each with 64 cores.. These estimations were performed using the Bilby package, employing the same waveform generator used to simulate our data. The results are presented in Table 1. The values reported in this table are obtained from those of Table 3 in Appendix A which shows all timing estimates for the entire set of waveforms. We observe that the CVAE is significantly more efficient than Bilby, achieving speeds about 4 orders of magnitude faster (occasionally even 5 orders of magnitude; see Table 3). To accelerate Bilby’s convergence, we use the true values as priors for all parameters except $y$ and $t_{M}$. For $y$ and $t_{M}$, we set uniform priors ranging from 0 to 3 and from 0 to 0.2 seconds, respectively. When the CVAE is used to provide priors for Bilby, the priors for these two lensing parameters are uniform distributions defined by the 95% confidence intervals of the CVAE’s posterior estimates, i.e. $\mu\pm 2\sigma$, where $\mu$ and $\sigma$ represent the posterior mean and standard deviation respectively. In this setup, the average runtime is reduced by 47.9% compared to the previous configuration.
@@
Figure 8: Box plot comparing KS statistics for uninformed/uninformed runs (blue) and CVAE-informed/uninformed runs (orange) for $t_{M}$ (top) and $y$ (bottom). The black horizontal lines inside the boxes represent the median value, the boxes represent the interquartile range (IQR) of the data. The whiskers extend up to 1.5 times the IQR, with values beyond this range considered outliers, shown as individual circles
In addition to significantly speeding up the inference, the use of the CVAE combined with Bilby does not degrade the accuracy of the parameter estimation obtained by Bilby when used in standalone mode. To demonstrate this, we performed 10 parameter estimation runs for both prior configurations on 20 randomly selected waveforms (indexed 1 to 20). Four of these waveforms (4, 16, 18, and 19) failed to converge in the uninformed runs and were excluded from the analysis; however, we note that the corresponding CVAE-informed runs did converge. For each remaining waveform, we computed the KS statistic between posteriors from different uninformed runs (excluding diagonal elements) and compared it to the KS statistic between uninformed and CVAE-informed runs. If the CVAE-informed prior leads to posterior distributions equivalent to those obtained with an uninformed prior, the KS statistic for CVAE-informed vs. uninformed runs should be comparable to that of different uninformed runs for the same waveform. The results, summarized in the box plots in Fig. 8, largely support this hypothesis: the median KS values for both comparisons lie within one quartile of each other, and outlier values, when they exist, appear in both cases. The only notable exception is waveform 9, where the KS statistic is significantly higher due to the CVAE-informed prior truncating a low-probability tail in the posterior. However, this does not affect the bulk of the distribution. Waveforms 10 and 11 show a large spread in KS statistic values due to noise-induced degeneracy in the matched-filter likelihood. This causes multiple convergence points in the ($tM$,$y$) plane for PE runs on waveforms 10 and 11 which leads to large KS statistics when comparing runs. The use of the CVAE prior restricts the sampling space and can mitigate this effect, as seen in waveform 11.